OpenStudy (anonymous):

It was reported that about 80% of airline tickets-nearly $65 billion worth last year are issued over the internet. What were the total sales of airline tickets last year? Round to the nearest billion.

OpenStudy (anonymous):

Well we know that 80% of some number, lets call it x, is 65 b. This means that x * 80% = 65 b Divide both sides by 80% x = 65 b / 80% = 65 b / 0.8 = 81.25 b Rounded to 81 Billions